OpenStudy (anonymous):

Find the sum of the first 20 terms of the geometric sequence {2,6,18,54,…}.

OpenStudy (anonymous):

Firstly find here r by dividing second term by first term..

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Yep.. So, here n = 20 you are given with: Use Sum formula: \[Sum = \frac{a(r^n - 1)}{r - 1}\]

OpenStudy (anonymous):

so the first term a = 2, ratio r = 3, number of terms n = 20 We know that sum of a GP Sn = a(r^n -1)/r-1 => 2(3^20-1)/3-1 => 2(3^20-1)/2 => 3^20 - 1 which comes out to be 3486784401 -1 = 3486784400.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You are welcome dear..